OpenStudy (anonymous):

u^2+4u+? to make equation a perfect square

OpenStudy (kainui):

What would the equation of a perfect square look like? Make up your own variable and then multiply it out. Then match up the corresponding parts to solve. =)

OpenStudy (neer2890):

?=4 u^2+4u+4 =(u+2)^2

OpenStudy (imstuck):

Do you know how to complete the square? This is what you would do to find that missing term. Do you know the process?

OpenStudy (imstuck):

neer2890 is correct, but maybe you should have been allowed to try it out yourself without being given the answer. Just sayin'.

OpenStudy (kainui):

I'll elaborate a little on my technique. You want a perfect square, then let's make one. \[\large (u+a)^2\] So what's our a? Well after we multiply it out we just get: \[\large u^2+2au+a^2 \] And we're saying that this is what our original problem is equal to. \[\Large u^2+2au+a^2 =u^2+4u+?\] So it looks like our question mark is whatever a^2 is. How do we solve for a? Well match up the terms and you can see that the middle terms match to give us this: \[\Large 2au=4u\]solve for a and then square it. That's what you must add.